Write the expression for the particular solution in integral form and solve it where possible.

y''+4y=f(x)
Answer #1

Som- criven, Y + uy = f(n) f(n) dua tuy ozy tuy = f(n) (32+4) -4 f(n) A.E, 2²+4=0 = 0² = -4 = 0 = £21 cf = e(e, cosant casin